A company uses Kshs. 10,000 worth of an item during the year.
T
he
ordering costs
are Kshs. 25 per
order and carrying charges
are 12.5% of th
e average inventory value. Find:
i.
T
he economic order quantity,
ii.
Number of orders per
year
iii.
Time period per order
iv.
The total cost.
1
Expert's answer
2018-11-08T11:26:09-0500

A company uses CR = Kshs. 10,000 worth of an item during the year. The ordering costs are C = Kshs. 25 per order and carrying charges are I = 12.5% = 0.125 of the average inventory value. Find:

i. The economic order quantity is EOQ = (2C*CR/I)^0.5 = (2*25*10,000/0.125)^0.5 = 2,000.

ii. Number of orders per year is n = CR/EOQ = 10,000/2,000 = 5.

iii. Time period per order is t = 1/n = 1/5 = 0.2 year = 73 days.

iv. The total cost is TC = CR + VC = CR + (2C*I*CR)^0.5 = 10,000 + (2*25*0.125*10,000)^0.5 = 10,000 + 250 = 10,250.